# DreamAnywhere: Object-Centric Panoramic 3D Scene Generation

### v1

@@ 3.2.1. Reference Image Generation
We synthesize a high-resolution reference image for each object by conditioning a diffusion model on the available multimodal cues. First, we extract a textual description of the object by prompting a vision-language model with a perspective view of a crop containing the object. This includes a description of the object’s visual attributes such as pose and color, serving as the primary conditioning signal. Second, we compute a partial depth map by combining the object’s segmentation mask with the estimated depth, offering geometric guidance for shape. Finally, we include a style image, reused from the panorama generation process, to ensure visual consistency and thematic coherence across the scene. As shown in Figure 5, our approach results in realistic and style-consistent object views. This approach is better suited for downstream 3D reconstruction compared to other diffusion-based amodal completion methods (see Figure 6).
3.2.2. Image-based Object Generation
Our object reconstruction pipeline first generates six multi-view images at fixed camera poses, and then transforms them into a NeRF representation that captures both the geometry and color of the object. We found NeRF to be more robust than 3DGS as a differentiable representation when learning the reconstruction. Therefore, we opt to transform it into 3DGS in a separate post-processing step. We leverage Zero123++ (Shi et al., 2023) for the multi-view generation task and use the triplane decoder of InstantMesh (Xu et al., 2024) to convert the multi-view images into NeRF.
@@ 3.2.2. Image-based Object Generation
In practice, generative models often introduce artifacts, including transparency, incomplete regions or structural inconsistencies. To enable a reliable and automated object reconstruction pipeline, we introduce a single-round refinement stage after generation. During this stage, we generate multiple object candidates and rank them by visual quality and alignment with the original input image using a VLM, similar to (Zhou et al., 2025). The VLM selects the best result from five candidates, one of which is generated using only the original cropped object image.
@@
Figure 5. Qualitative evaluation of our object reconstruction module. Employing multimodal cues faithfully captures the stylistic and structural essence of low-fidelity or incomplete objects, reconstructing a suitable 3D counterpart to be later placed in the environment.
Figure 6. Diffusion-based amodal completion methods (such as pix2gestalt (Ozguroglu et al., 2024) and Gen3DSR (Dogaru et al., 2024)) face challenges in producing high-quality inputs for multi-view image generation. Our method prioritizes structural fidelity by resynthesizing the reference image to yield higher quality 3D reconstructions.
3.2.3. Pose Estimation
@@
We decompose object pose estimation into two parts: the absolute pose of the original object $O_{\text{crop}}$ with respect to the viewing origin and the relative transformation between $O_{\text{crop}}$ and its high-quality counterpart $O_{\text{hq}}$. The absolute pose is estimated by projecting the crop’s midpoint along a ray from the camera origin to the estimated mean depth. Orientation and scale are heuristically derived based on the crop geometry and panorama alignment. This provides a coarse but coherent placement aligned with the image-plane and panorama geometry. To semantically align the point cloud of the high-quality object $O_{\text{hq}}$ with the original $O_{\text{crop}}$ we compute the relative pose between them. We employ MAST3R (Leroy et al., 2024) on image pairs (perspective projections) of the original and high-quality object to obtain the relative pose. The final pose of $O_{\text{hq}}$ preserves semantic alignment with the original object’s placement.
3.3. 3D Background Generation
@@
Generating a 3D scene suitable for exploration from a single 360° image is an underconstrained problem, as the available information is limited to a single viewpoint. To address this, we employ a three-step process that reconstructs missing scene content through a combination of inpainting and score distillation sampling (SDS) (Poole et al., 2023). First, we optimally initialize the 3DGS point cloud from the panoramic image and prepare disocclusions for inpainting through a pretuning stage; Second, we incrementally populate disocclusions with new Gaussians through a 3D inpainting process conditioned on the existing point cloud; Finally, we ensure multi-view consistency by applying SDS on a larger set of possible views. During all optimization steps, we leverage MCMC-densification (Kheradmand et al., 2024), a process demonstrated in Figure 7.
3.3.1. Background Initialization and Pretuning.
@@
We initialize the 3DGS point cloud by unprojecting the pixels through a spherical camera placed at the origin, where the covariance matrices are initialized from normals derived from the depth map $D_{B}$, resulting in pixel-tight views when viewed from the origin. For pretuning, we optimize the initial Gaussians using mixed supervision from an upscaled version of the background panorama $I_{B}$ (90%) and random views (10%). The $I_{B}$ loss follows standard 3DGS supervision, while the random-view supervision focuses on identifying dissocluded regions.
@@
Figure 7. The background generation stage starts by identifying large occluded regions and preparing the 3DGS point cloud for inpainting, then proceeds to incrementally fill in new content and finally enforces multi-view consistency when rendering from novel views with score distillation.
3.3.2. Incremental Inpainting
@@
To synthesize realistic new content in disoccluded regions, we leverage a pre-trained inpainting model conditioned on the rendered projections of the existing 3DGS splats. Specifically, we select a set of keyframes that target important regions with missing content. For each keyframe we render the point cloud and generate an inpainted RGB image and an inpainting mask, computed by applying max-pooling (kernel size 25) and thresholding ($T{<}0.5$) on the transmittance map $T$, isolating low-opacity regions as targets for inpainting. For each pixel in the target region, we instantiate a new Gaussian by unprojecting it with a scene-aligned depth estimate. This depth is obtained using a diffusion model repurposed and fine-tuned for 3DGS depth inpainting (Liu et al., 2024), and provides a strong geometric prior. The inpainting process is done incrementally per keyframe, minimizing inconsistencies across overlapping views, similarly to (Höllein et al., 2023).
3.3.3. Multi-view Fine-tuning
@@
Finally, to consolidate the inpainted views with the existing 3D content, we employ mixed supervision from the panorama image (25%), the inpainted images (25%) and SDS (Poole et al., 2023) from random views around the inpainted regions (50%). This final stage of SDS fine-tuning helps to remove view-inconsistencies and ensures that the Gaussians introduced during inpainting are consistent and aligned with the existing scene.
3.4. Composition
@@
To further improve visual coherence, we address two key sources of error that can disrupt immersion. The first is the elimination of contact shadows — an unavoidable consequence of the mask dilation step used during object removal. Since achieving full control over lighting in diffusion models (Guo et al., 2023; Corvi et al., 2023) and relighting remain challenging problems, we reintroduce shadows through a splat-aware shadow mapping pass (details in the supplemental). The second issue involves inaccurate pose estimation, which can lead to objects violating physical plausibility. To mitigate this, we automatically snap object positions to dominant support planes (e.g. floor or walls), ensuring a more realistic and physically grounded scene integration.
4. Results
In this section we provide a quantitative and qualitative evaluation on the scene generation task, as well as an ablation of the core components of our pipeline. To assess the effectiveness of our approach in generating 3D scenes from text prompts, we focus on its ability to produce coherent and immersive scenes with high-quality and structurally consistent novel view synthesis. We compare our method against three state-of-the-art approaches for immersive scene generation: LayerPano3D (Yang et al., 2024c), Text2Room (Höllein et al., 2023), and DreamScene360 (Zhou et al., 2025). Additionally, in Figure 8 we provide a qualitative comparison with other methods that share a similar goal but for which a direct comparison is non-trivial.
Our method is evaluated on 17 text prompts commonly found in recent literature (Shriram et al., 2024; Yang et al., 2024c). For each scene we recorded a 6-second walk-through video (at $1024\times 1024$px resolution) along a predefined camera trajectory different from the one used for the optimization described in Section 3.3, covering large parts of the scene’s geometry as well as occluded surfaces. Examples of video frames are shown in Figure 11.
We generated all scenes on a single NVIDIA H100, with each scene typically taking 10-15m to generate, given an average of 7.5 objects per scene. More results can be found in Figure 12.
@@ 4. Results
Figure 8. Qualitative evaluation against WonderWorld (Yu et al., 2024a), DreamScene (Li et al., 2025) and ViewCrafter (Yu et al., 2024c). WonderWorld degenerates into flat, view-dependent renderings when shown from out-of-context viewpoints, whereas DreamScene exhibits reduced fidelity in complex regions and object poses while expecting the positioning of objects to be known a priori. ViewCrafter combines video diffusion and reconstruction models but it is not able to maintain 3D coherence for long camera paths.
Figure 9. Qualitative ablation of out-of-domain sampling for panorama generator models. Our method captures the stylistic guidance faithfully, but without IP-Adapter it reverts back to its base panoramic distribution similarly to other recent methods. Simply using a perspective IP-Adapter does not work because of the mismatching distributions at training time.
Figure 10. User study results for pairwise comparisons between our method, DreamScene360, Text2Room and LayerPano3D. Ratings are given on a 5-point Likert scale capturing strong, weak and no preference between the approaches. The participants rated our approach higher than the comparison methods on all axes.
4.1. Quantitative Evaluation
Table 1. Quantitative evaluation of our method against Text2Room (Höllein et al., 2023), DreamScene360 (Zhou et al., 2025) and Layerpano3D (Yang et al., 2024c). Metrics are averaged across all scenes and computed on 136 frames per scene. Our method has state-of-the-art performance in terms of perceptual image quality and visual aesthetics. The differences in CLIP scores (evaluated against the original text prompt) are small and should be interpreted with care (see main text).
@@ 4.1. Quantitative Evaluation
We quantitatively evaluate prompt alignment and image quality. To measure image quality and aesthetics, we use CLIP-IQA+ (Wang et al., 2023), Q-Align and A-Align (Wu et al., 2023a) metrics. To quantify prompt alignment we compute the CLIPScore (Hessel et al., 2021) between the prompt and each rendered frame. The results in Table 1 demonstrate that our method outperforms competing work in terms of both image quality and aesthetics metrics. For completeness, we also evaluate CLIP scores on single views against the scene’s text prompt. However, individual views typically capture only a small portion of the scene, making CLIP scores a less reliable metric for our novel view evaluation.
4.2. Qualitative Evaluation
4.2.1. Comparison
@@
4.2.2. Ablation
We present an ablation of the image generator architecture and the necessity for the joint fine-tuning procedure in Figure 9, comparing it to (Yang et al., 2024c) and (Zhang et al., 2024). Despite training on only a subset of the public data used by (Yang et al., 2024c), we demonstrate stronger prompt adherence.
@@ 4.2.2. Ablation
Our object generation module effectively addresses several common challenges in object extraction from images, such as panoramic distortion, low-resolution crops, and imperfect segmentation masks. Examples of these issues are illustrated in Figure 6. One alternative is to apply super-resolution and amodal completion but this approach typically requires multiple specialized modules, each with its own issues (Figure 5). Generating the same level of detail with super-resolution techniques on the panoramic image is still a challenge and requires much larger network architectures.
4.3. User Study
[…]
5. Discussion and Limitations
The modular nature of our pipeline has the disadvantage of having multiple possible points of failure. Conversely, this very modularity, serves as a significant advantage for targeted user intervention, which is often desired in generative applications. It also facilitates easy replacement of individual modules (e.g. the image generator method with (Yang et al., 2024c) or the 3D inpainting step with (Yu et al., 2024c)). A core strength of our method lies in its ability to generate globally coherent 3D scenes that remain consistent across a wide range of viewpoints. While this comes at the cost of localized sharpness compared to models optimized for narrow view paths, it enables free exploration without structural artifacts or semantic collapse. In contrast, video diffusion-based approaches typically reconstruct scenes from a predefined camera trajectory, yielding sharper results along that path but failing to generalize beyond it. This makes direct comparisons inherently asymmetric. To provide a fairer assessment, we demonstrate in Figure 12 and the accompanying video how our method can optionally be combined with a video-to-video diffusion model (e.g., Cosmos-Transfer1 (Alhaija et al., 2025)) to enhance rendering sharpness along known view paths. This hybrid configuration showcases the flexibility of our modular pipeline and illustrates the upper bound of visual quality achievable with targeted post-processing, making our system competitive even in restricted-view settings.
6. Conclusion

### v2

@@ 3.2.1 Reference Image Generation
We synthesize a high-resolution reference image for each object by conditioning a diffusion model on the available multimodal cues. First, we extract a textual description of the object by prompting a vision-language model with a perspective view of a crop containing the object. This includes a description of the object’s visual attributes such as pose and color, serving as the primary conditioning signal. Second, we compute a partial depth map by combining the object’s segmentation mask with the estimated depth, offering geometric guidance for shape. Finally, we include a style image, reused from the panorama generation process, to ensure visual consistency and thematic coherence across the scene. As shown in Figure 5, our approach results in realistic and style-consistent object views. This approach is better suited for 3D reconstruction compared to other diffusion-based amodal completion methods (see Figure 11).
3.2.2 Image-based Object Generation
@@
Our object reconstruction pipeline first generates six multi-view images at fixed camera poses, and then transforms them into a NeRF representation that captures both the geometry and color of the object. We found NeRF to be more robust than 3DGS as a differentiable representation when learning the reconstruction. Therefore, we opt to transform it into 3DGS in a separate post-processing step. We leverage Zero123++ [39] for the multi-view generation task and use the triplane decoder of InstantMesh [52] to convert the multi-view images into NeRF. In practice, generative models often introduce artifacts, including transparency, incomplete regions or structural inconsistencies. To enable a reliable and automated object reconstruction pipeline, we introduce a single-round refinement stage after generation. During this stage, we generate multiple object candidates and rank them by visual quality and alignment with the original input image using a VLM, similar to [71]. The VLM selects the best result from five candidates, one of which is generated using only the original cropped object image.
@@
Figure 5: Qualitative evaluation of our object reconstruction module. Employing multimodal cues faithfully captures the stylistic and structural essence of low-fidelity or incomplete objects, reconstructing a suitable 3D counterpart to be later placed in the environment.
3.2.3 Pose Estimation
@@
We decompose object pose estimation into two parts: the absolute pose of the original object $O_{\text{crop}}$ with respect to the viewing origin and the relative transformation between $O_{\text{crop}}$ and its high-quality counterpart $O_{\text{hq}}$. The absolute pose is estimated by projecting the crop’s midpoint along a ray from the camera origin to the estimated mean depth. Orientation and scale are heuristically derived based on the crop geometry and panorama alignment. This provides a coarse but coherent placement aligned with the image-plane and panorama geometry. To semantically align the point cloud of the high-quality object $O_{\text{hq}}$ with the original $O_{\text{crop}}$ we compute the relative pose between them. We employ MAST3R [23] on image pairs (perspective projections) of the original and high-quality object to obtain the relative pose. The final pose of $O_{\text{hq}}$ preserves semantic alignment with the original object’s placement.
3.3 3D Background Generation
@@
Generating a 3D scene suitable for exploration from a single 360° image is an underconstrained problem, as the available information is limited to a single viewpoint. To address this, we employ a three-step process that reconstructs missing scene content through a combination of inpainting and score distillation sampling (SDS) [34]. First, we optimally initialize the 3DGS point cloud from the panoramic image and prepare disocclusions for inpainting through a pretuning stage; Second, we incrementally populate disocclusions with new Gaussians through a 3D inpainting process conditioned on the existing point cloud; Finally, we ensure multi-view consistency by applying SDS on a larger set of possible views. During all optimization steps, we leverage MCMC-densification [22], a process demonstrated in Figure 6.
3.3.1 Background Initialization and Pretuning.
@@
We initialize the 3DGS point cloud by unprojecting the pixels through a spherical camera placed at the origin, where the covariance matrices are initialized from normals derived from the depth map $D_{B}$, resulting in pixel-tight views when viewed from the origin. For pretuning, we optimize the initial Gaussians using mixed supervision from an upscaled version of the background panorama $I_{B}$ (90%) and random views (10%). The $I_{B}$ loss follows standard 3DGS supervision, while the random-view supervision focuses on identifying disoccluded regions.
@@
Figure 6: The background generation stage starts by identifying large occluded regions and preparing the 3DGS point cloud for inpainting, then proceeds to incrementally fill in new content and finally enforces multi-view consistency when rendering from novel views with score distillation.
3.3.2 Incremental Inpainting
@@
To synthesize realistic new content in disoccluded regions, we leverage a pre-trained inpainting model conditioned on the rendered projections of the existing 3DGS splats. Specifically, we select a set of keyframes that target important regions with missing content. For each keyframe we render the point cloud and generate an inpainted RGB image and an inpainting mask, computed by applying max-pooling (kernel size 25) and thresholding (T<0.5T{<}0.5) on the transmittance map $T$, isolating low-opacity regions as targets for inpainting. For each pixel in the target region, we instantiate a new Gaussian by unprojecting it with a scene-aligned depth estimate. This depth is obtained using a diffusion model repurposed and fine-tuned for 3DGS depth inpainting [28], and provides a strong geometric prior. The inpainting process is done incrementally per keyframe, minimizing inconsistencies across overlapping views, similarly to [18].
3.3.3 Multi-view Fine-tuning
@@
Finally, to consolidate the inpainted views with the existing 3D content, we employ mixed supervision from the panorama image (25%), the inpainted images (25%) and SDS [34] from random views around the inpainted regions (50%). This final stage of SDS fine-tuning helps to remove view-inconsistencies and ensures that the Gaussians introduced during inpainting are consistent and aligned with the existing scene as shown in Figure 6.
3.4 Composition
@@
To further improve visual coherence, we address two key sources of error that can disrupt immersion. The first is the elimination of contact shadows—an unavoidable consequence of the mask dilation step used during object removal. Since achieving full control over lighting in diffusion models [15, 10] and relighting remain challenging problems, we reintroduce shadows through a splat-aware shadow mapping pass (details in the supplemental). The second issue involves inaccurate pose estimation, which can lead to objects violating physical plausibility. To mitigate this, we automatically snap object positions to dominant support planes (e.g. floor or walls), ensuring a more realistic and physically grounded scene integration.
4 Results
In this section we provide a quantitative and qualitative evaluation on the scene generation task, as well as an ablation of the core components of our pipeline. To assess the effectiveness of our approach in generating 3D scenes from text prompts, we focus on its ability to produce coherent and immersive scenes with high-quality and structurally consistent novel view synthesis. We compare our method against three state-of-the-art approaches for immersive scene generation: LayerPano3D [53], Text2Room [18], and DreamScene360 [71]. Additionally, in Figure 7 we provide a qualitative comparison with other methods that share a similar goal but would require substantial integration effort to directly compare against.
Our method is evaluated on 17 text prompts commonly found in recent literature [40, 53]. For each scene we recorded a 6-second walk-through video (at $1024\times 1024$px resolution) along a predefined camera trajectory different from the one used for the optimization described in Section 3.3, covering large parts of the scene’s geometry as well as occluded surfaces. Examples of video frames are shown in Figure 10. We generated all scenes on a single H100, with each scene taking around 15 mins to generate, given an average of 7.5 objects per scene. We include more results in the supplementary material.
@@ 4 Results
Figure 7: Qualitative evaluation against WonderWorld [58], DreamScene [25] and ViewCrafter [60]. WonderWorld degenerates into flat, view-dependent renderings when shown from out-of-context viewpoints, whereas DreamScene exhibits reduced fidelity in complex regions and object poses while expecting the positioning of objects to be known a priori. ViewCrafter combines video diffusion and reconstruction models but it is not able to maintain 3D coherence for long camera paths.
Figure 8: Qualitative ablation of out-of-domain sampling for panorama generator models. Our method captures the stylistic guidance faithfully, but without IP-Adapter it reverts back to its base panoramic distribution similarly to other recent methods. Simply using a perspective IP-Adapter does not work because of the mismatching distributions at training time.
Figure 9: User study results for pairwise comparisons between our method, DreamScene360, Text2Room and LayerPano3D. Ratings are given on a 5-point Likert scale capturing strong, weak and no preference between the approaches. The participants rated our approach higher than the comparison methods on all axes.
4.1 Quantitative Evaluation
Table 1: Quantitative evaluation of our method against Text2Room [18], DreamScene360 [71] and Layerpano3D [53]. Metrics are averaged across all scenes and computed on 136 frames per scene. Our method has state-of-the-art performance in terms of perceptual image quality and visual aesthetics. The differences in CLIP scores (evaluated against the original text prompt) are small and should be interpreted with care (see main text).
@@ 4.1 Quantitative Evaluation
Figure 10: Qualitative comparison with Text2Room [18], DreamScene360 [71] and LayerPano3D [53]. Whereas scenes from these competing method suffer from disturbing artifacts for significant viewpoint changes, our method achieves coherent novel view synthesis for virtually all viewpoints.
Figure 11: Diffusion-based amodal completion methods (such as pix2gestalt [30]) face challenges in producing high-quality inputs for multi-view image generation. Our method prioritizes structural fidelity by resynthesizing the reference image to yield higher quality 3D reconstructions.
Table 2: Ablation of the perspective conditioned IP-adapter for panorama generation. Out-of-domain prompt following improves without compromising generation quality. FID, KID and IS are evaluated on the test split of Matterport3d [4]. The relative difference in distribution similarity is significantly less relevant than CLIP improvement.
[…]
Full model 26.49 25.18 73.09 2.04 7.51
w/o IP-adapter 24.82 24.10 74.34 2.40 7.14
w/o joint fine-tuning 17.54 14.48 437.67 33.91 2.98
We quantitatively evaluate prompt alignment and image quality. To measure image quality and aesthetics, we use CLIP-IQA+ [48], Q-Align and A-Align [49] metrics. To quantify prompt alignment we compute the CLIPScore [17] between the prompt and each rendered frame. The results in Table 1 demonstrate that our method outperforms competing work in terms of both image quality and aesthetics metrics. For completeness, we also evaluate CLIP scores on single views against the scene’s text prompt. However, individual views typically capture only a small portion of the scene, making CLIP scores a less reliable metric for our novel view evaluation.
@@
4.2 User Study
To assess the effectiveness of our model in generating immersive and navigable 3D scenes, we conducted a user study with 28 participants, including graduate students and professionals in graphics and vision. Participants evaluated short videos of scenes on: Coherence: Do the spatial layout, object placement, and interactions in the scene form a logical and believable whole?; Compatibility: Do the objects, materials, lighting, and textures align well with the given text prompt?; Immersiveness: Does the scene convincingly represent a 3D environment, evoking a sense of depth, realism and presence?; Overall Preference. Each comparison was rated from 2-2 to +2+2. As shown in Figure 9, our method was consistently preferred across all metrics and in particular was praised for its geometric structure in novel view synthesis, as shown by the coherence and immersiveness scores. The results were statistically significant ($p<.001$) according to the Wilcoxon signed-rank tests.
4.2.1 Ablation
We present a quantitative ablation of the panorama generator in Table 2, highlighting the need for joint fine-tuning as well as the substantial prompt adherence obtained with the introduction of the IP-Adapter mechanism. A qualitative comparison with [53] and [63] is shown in Figure 8, noting that each model is built on top of different foundation models. Despite relying on only a subset of the public data used by [53], our approach allows for more style-adhering images.
@@ 4.2.1 Ablation
Our object generation module effectively addresses several common challenges in object extraction from images, such as panoramic distortion, low-resolution crops, and imperfect segmentation masks. Examples of these issues are illustrated in Figure 11. One alternative is to apply super-resolution and amodal completion but this approach typically requires multiple specialized modules, each with its own issues (Figure 5). Generating the same level of detail with super-resolution techniques on the panoramic image is still a challenge.
5 Conclusion